OpenStudy (jenniferjuice):

In a certain county, the number of Charter Schools is eleven less than twice the number of Alternative Schools. There are 35 Charter Schools in the county. Which of the following equations would you use to find the number of Alternative Schools, a, in the county?

OpenStudy (anonymous):

c = 2a - 11 c = 35 .'. 35 = 2a - 11 -> 35 + 11 = 2a -> 46 = 2a -> 46/2 = a -> 23 = a 35 = 2(23) - 11 46 = 2(23) 46 = 46 qed

OpenStudy (jenniferjuice):

2a − 11 = 35 2a = 35 − 11 2(a − 11) = 35 Quantity of a minus 11 all over 2 equals 35

OpenStudy (jenniferjuice):

^^^options^^^

OpenStudy (anonymous):

Clearly the first is correct, then.

OpenStudy (jenniferjuice):

alrighty thankyou